# Generalized discrete Markov spectra

### v2

@@ 1.1. Background
Theorem 1.1 (Markov’s theorem [mar1, mar2]).
@@
This is studied in the context of Diophantine approximation, which concerns approximating irrational numbers by rationals. Hurwitz [hurwitz] pointed out that the subset $\mathcal{L}_{<3}$ of the Markov-Lagrange spectrum below 3 coincides with the discrete Markov spectrum $\mathcal{M}_{d}$, that is,
@@
(see [cusick-flahive]). For further details of the history, see [aig, cusick-flahive, reutenauer].
@@ 1.2. Main Results
where $k_{1},k_{2},k_{3}$ are nonnegative integers. A positive integer appearing in a solution to this equation is called a $(k_{1},k_{2},k_{3})$-generalized Markov number, or $(k_{1},k_{2},k_{3})$-GM number for short. This equation was introduced by [gyo21, bana] for the case $k_{1}=k_{2}=k_{3}=1$ and by [gyomatsu] for the general case in the context of generalized cluster algebras. It has been studied from various perspectives as an equation possessing properties similar to those of the Markov equation (for example, [banaian, banaian-sen, bana-gyo, chen-li, courcy-litman-mizuno, gyoda-maruyama-sato, chen-li2, bao-li, chen-jia]).
@@
Clearly, we have $\mathcal{M}_{d}=\mathcal{M}_{0,0,0}$. Therefore we call $\mathcal{M}_{k_{1},k_{2},k_{3}}$ the $(k_{1},k_{2},k_{3})$-generalized discrete Markov spectrum. In this paper, we compute such $\alpha$ using loops on the once-punctured torus together with their associated periodic continued fractions, and we express the corresponding (generalized) Markov numbers as the numbers of perfect matchings of graphs called snake graphs. A crucial feature of this approach is that it does not rely on Christoffel words (also called Cohn words), which are often used in modern proofs of the Markov theorem, such as those by [aig], [bombieri], and [reutenauer]. In these works, both the Cohn matrices and the strongly admissible sequences, which are indispensable tools to prove Markov’s theorem, are constructed from Christoffel words. This approach simplifies the argument and has thus become the standard method. However, in the generalized setting discussed in this paper, neither the generalized Cohn matrices nor the generalized strongly admissible sequences can be constructed from Christoffel words. Consequently, the proofs that rely on Christoffel words cannot be extended in a straightforward way. In this respect, the snake graphs used in this paper retain more detailed information than Christoffel words and are appropriate to use in the proof of the generalized theorem. We would like to emphasize that, in addition to introducing a generalization of the Markov spectrum, this paper also aims to propose a new proof technique for the classical Markov theorem.
@@
We further establish several important properties of the generalized discrete Markov spectra. The first concerns the relationship between the classical discrete Markov spectrum and the $(2,2,2)$-generalized discrete Markov spectrum. It was already known, by [gyomatsu], that classical Markov numbers and $(2,2,2)$-GM numbers are connected through a remarkable relation involving squaring and taking square roots. We now show that a similarly simple and elegant relation holds between their spectra.
@@
However, this question has already been answered in the negative by [gyomatsu]. Therefore, based on the main results, we propose a generalization of the equivalent form of the uniqueness conjecture, for which no counterexamples are currently known (Conjecture 7.6).
@@ 2.1. Markov Numbers and Markov Tree
Proposition 2.1 (see [gyoda-maruyama-sato]*Proposition 3.2).
@@
Lemma 2.6 is essentially proved in [gyoda-maruyama-sato]*Proposition 7.22 (2) (the case $k=0$). Indeed, if $(r,t,s)\in\mathrm{F}\mathbb{T}$, then $(\frac{1}{s},\frac{1}{t},\frac{1}{r})$ appears in $\mathrm{F}\mathbb{T}$ at the position symmetric to $(r,t,s)$ with respect to the central axis of the tree. Moreover, since $(n_{\frac{1}{s}},n_{\frac{1}{t}},n_{\frac{1}{r}})=(n_{s},n_{t},n_{r})$, the element $v_{t}^{+}$ in [gyoda-maruyama-sato] coincides with $u_{\frac{1}{t}}$.
@@ 2.2. Cohn Words, Cohn Matrices, and Cohn Tree
The Cohn matrix defined here differs from that in [aig] in the arrangement of its entries. More precisely, the present definition is obtained from the one in [aig] by swapping the diagonal entries.
@@
A proof of Theorem 2.11 can be found in [aig]*Theorem 4.13. In [aig], only the case $t\in[0,1]$ is treated, but the same argument applies to $t\in(1,\infty]$ as well.
@@ 2.3. Continued Fraction Matrix Decompositions of Cohn Matrices
Next, we define the strongly admissible sequence s(t)s(t) [^1] In [aig], the term “strongly admissible sequence” is used for the bi-infinite sequence $^{\infty}s(t)^{\infty}$ in the sense of this paper, whereas here we use this name for its single block $s(t)$. . For $t=\frac{0}{1}$ or $t=\frac{1}{0}$, set $s(\frac{0}{1})=s(\frac{1}{0})=(1,1)$. For $t\in(0,\infty)$, we define $s(t)$ as follows:
@@
    If $t\in[1,\infty]$, then $c_{t}$ consists of letters $q$ and $r$, and $s(t)$ is obtained from $c_{t}$ by replacing each $q$ with “2,22,2” and each $r$ with “1,11,1”. In [aig, bombieri], $s(t)$ is constructed from $c_{t}$ by using this replacement.
@@
By Remark 2.15 (2), $s(\frac{1}{t})$ is obtained from $s(t)$ by replacing the initial $2,2$ with $2,1,1$ and the final $1,1$ with $2$. The lemma follows from direct matrix computation (see also [bombieri2]*Corollary 24). ∎
@@ 2.4. Markov-Lagrange Spectrum and Markov Spectrum
We have the following theorem [^2] In [hurwitz], a more general formula is given for computing $\mathcal{L}(\alpha)$ for an arbitrary irrational number $\alpha$ using its continued fraction expansion and $\lim\sup$. By applying this formula to the case of quadratic irrationals, one obtains Theorem 2.20 stated in this section. For a detailed discussion of this reformulation, see [bombieri]. :
Theorem 2.20 ([hurwitz]).
@@
The quadratic form $Q_{T}$ has rational coefficients. We have the following [^3] In [perron2], a more general formula is given for computing $\mathcal{M}(Q)$ for arbitrary irreducible quadratic form $Q$ using its continued fraction expansion and $\sup$. By applying this formula to the case of quadratic irrationals, one obtains Theorem 2.22 stated in this section. For a detailed discussion of this reformulation, see [classical-dynamical]*Appendix C.:
Theorem 2.22 ([perron2]).
@@ 3. Generalized Markov Equation and Generalized Markov Numbers
${((1,1),(4,2),(1,3))}$${((4,2),(21,1),(1,3))}$${((1,1),(17,3),(4,2))}$${((21,1),(121,2),(1,3))}$${((4,2),(457,3),(21,1))}$${((17,3),(373,1),(4,2))}$${((1,1),(81,2),(17,3))}$${((121,2),(703,1),(1,3))\cdots}$${((21,1),(15082,3),(121,2))\cdots}$${((457,0),(57121,2),(21,1))\cdots}$${((4,2),(10033,1),(457,3))\cdots}$${((373,1),(8185,3),(4,2))\cdots}$${((17,3),(38025,2),(373,1))\cdots}$${((81,2),(8227,1),(17,3))\cdots}$${((1,1),(386,3),(81,2))\cdots}$
@@
It follows from [bana-gyo]*Proposition 3.8 and [gyomatsu]*Theorem 1. ∎
@@
    If $b=\max\{a,c\}$, then $(a,b,c)=(1,1,1)$ ([gyomatsu]*Lemma 4), therefore the set
@@
Let $(n^{\ast}_{\frac{1}{t}},i^{*}_{\frac{1}{t}})$ be a $(k_{1},k_{2},k_{3},\sigma^{\ast})$-GM number-position pair associated with $t$. Then, by Theorem 3.2, we have
@@
$u_{\frac{1}{t}}=n_{t}-u^{\ast}_{t}-k_{t}.$
The proof of Lemma 3.6 can be seen [gyoda-maruyama-sato]*Proposition 7.22 (2) for the case where $k_{1}=k_{2}=k_{3}$. We can prove the general case in the same way. As in the case of Lemma 2.6, replace $v_{t}^{+}$ in [gyoda-maruyama-sato] with $u^{*}_{\frac{1}{t}}$ and take the duality $(\ )^{\ast}$.
@@ 4.1. Generalized Cohn Matrices
In [gyo-maru]*Lemma 4.5, Theorem 4.2 is treated for $k_{1}=k_{2}=k_{3}$ and $t\in[0,1]$. The general case is proved similarly.
@@ 4.2. Continued Fraction Matrix Decomposition of Generalized Cohn Matrices
To prove Theorem 4.6, we introduce the snake graph. Assume $a_{i}\in\mathbb{Z}_{\geq 1}$. First, recall the snake graph associated with a continued fraction $a_{1},\dots,a_{\ell}]$ according to [cs18]:
@@
In [cs18], the authors establish the following relation between a continued fraction and its associated snake graph:
Theorem 4.11 ([cs18]*Theorem 3.4).
@@
If $a_{1},\dots,a_{n}]=\frac{p_{n}}{q_{n}}$, then the corresponding product of continued fraction matrices is given by (see [bombieri]*Section 2.1):
@@
On the other hand, [gyoda-maruyama-sato]*Theorem 7.10 provides an expression for $C_{t}$ for $t\in(0,1)$. Adapting it to the notation of the present paper, we obtain the following:
@@
Note that when comparing Theorem 4.12 with [gyoda-maruyama-sato]*Theorem 7.10 (2), the positions of the matrix entries differ due to the difference in the convention of Cohn matrices described in Remark 2.10.
@@
To prove Theorem 4.14, we need the snake graph calculus. The following two lemmas are essentially proved by [banaian-sen]. In this paper, the statement is formulated in terms of the numerators of continued fractions, but by applying the Theorem 4.11 it can be rephrased as a statement about the number of perfect matchings.
Lemma 4.16 ([banaian-sen]*Lemma 3).
@@
For the first statement of Lemma 4.17, the case where $k=1$ is proved in [banaian-sen]*Lemma 5. We can prove other cases in the same way as [banaian-sen]*Lemma 5. The second statement follows from Proposition 4.10. ∎
@@
The special case of Lemma 4.18 is proved by [gyoda-maruyama-sato]*(7.2). Lemma 4.18 can be proved in the same way.
@@ 5. Generalized Markov Length and Generalized Markov Distance
Although the curve $\gamma_{AB}$ depends on whether the interior of the line segment from $A$ to $B$ is shifted to the left or right, the value of $d(A,B)$ is independent of this choice. This fact is proved by [banaian]*Lemma 8 for $k_{1}=k_{2}=k_{3}$, and we can apply the same argument to general cases.
@@
Theorem 5.7 ([banaian]*Proposition 2 and Lemma 9).
@@
In [banaian]*Lemma 8, only the case $k_{1}=k_{2}=k_{3}$ is treated, but the theorem applies in the same way even when these three values are distinct. In the case $k_{1}=k_{2}=k_{3}=0$, Theorem 5.7 is proved by [llrs]*Theorem 3.5.
@@ 6.1. Main theorems and their Proofs
$\ell(^{\infty}S|S^{\infty})=\frac{\sqrt{\mathrm{tr}(CF_{S})^{2}-(-1)^{n}4}}{(CF_{S})_{21}},$
@@
$F_{t}=\left\{\frac{\sqrt{\mathrm{tr}(CF_{S})^{2}-4}}{(CF_{S})_{21}}\ \middle|\ S=(a_{k}\dots,a_{k+n-1})\text{ for some $k$ }\right\},$
@@
By general theory of continued fraction matrix (see [bombieri2]*Appendix A), we have

### v3

@@ 1.1. Background
Theorem 1.1 (Markov’s theorem [23, 24]).
@@
This is studied in the context of Diophantine approximation, which concerns approximating irrational numbers by rationals. Hurwitz [20] pointed out that the subset $\mathcal{L}_{<3}$ of the Markov-Lagrange spectrum below 3 coincides with the discrete Markov spectrum $\mathcal{M}_{d}$, that is,
@@
(see [13]). For further details of the history, see [1, 13, 27].
@@ 1.2. Main Results
where $k_{1},k_{2},k_{3}$ are nonnegative integers. A positive integer appearing in a solution to this equation is called a $(k_{1},k_{2},k_{3})$-generalized Markov number, or $(k_{1},k_{2},k_{3})$-GM number for short. This equation was introduced by [19, 4] for the case $k_{1}=k_{2}=k_{3}=1$ and by [18] for the general case in the context of generalized cluster algebras. It has been studied from various perspectives as an equation possessing properties similar to those of the Markov equation (for example, [5, 3, 2, 12, 14, 16, 11, 6, 10]).
@@
Clearly, we have $\mathcal{M}_{d}=\mathcal{M}_{0,0,0}$. Therefore we call $\mathcal{M}_{k_{1},k_{2},k_{3}}$ the $(k_{1},k_{2},k_{3})$-generalized discrete Markov spectrum. In this paper, we compute such $\alpha$ using loops on the once-punctured torus together with their associated periodic continued fractions, and we express the corresponding (generalized) Markov numbers as the numbers of perfect matchings of graphs called snake graphs. A crucial feature of this approach is that it does not rely on Christoffel words (also called Cohn words), which are often used in modern proofs of the Markov theorem, such as those by [1], [7], and [27]. In these works, both the Cohn matrices and the strongly admissible sequences, which are indispensable tools to prove Markov’s theorem, are constructed from Christoffel words. This approach simplifies the argument and has thus become the standard method. However, in the generalized setting discussed in this paper, neither the generalized Cohn matrices nor the generalized strongly admissible sequences can be constructed from Christoffel words. Consequently, the proofs that rely on Christoffel words cannot be extended in a straightforward way. In this respect, the snake graphs used in this paper retain more detailed information than Christoffel words and are appropriate to use in the proof of the generalized theorem. We would like to emphasize that, in addition to introducing a generalization of the Markov spectrum, this paper also aims to propose a new proof technique for the classical Markov theorem.
@@
We further establish several important properties of the generalized discrete Markov spectra. The first concerns the relationship between the classical discrete Markov spectrum and the $(2,2,2)$-generalized discrete Markov spectrum. It was already known, by [18], that classical Markov numbers and $(2,2,2)$-GM numbers are connected through a remarkable relation involving squaring and taking square roots. We now show that a similarly simple and elegant relation holds between their spectra.
@@
However, this question has already been answered in the negative by [18]. Therefore, based on the main results, we propose a generalization of the equivalent form of the uniqueness conjecture, for which no counterexamples are currently known (Conjecture 7.6).
@@ 2.1. Markov Numbers and Markov Tree
Proposition 2.1 (see [16]*Proposition 3.2).
@@
Lemma 2.6 is essentially proved in [16]*Proposition 7.22 (2) (the case $k=0$). Indeed, if $(r,t,s)\in\mathrm{F}\mathbb{T}$, then $(\frac{1}{s},\frac{1}{t},\frac{1}{r})$ appears in $\mathrm{F}\mathbb{T}$ at the position symmetric to $(r,t,s)$ with respect to the central axis of the tree. Moreover, since $(n_{\frac{1}{s}},n_{\frac{1}{t}},n_{\frac{1}{r}})=(n_{s},n_{t},n_{r})$, the element $v_{t}^{+}$ in [16] coincides with $u_{\frac{1}{t}}$.
@@ 2.2. Cohn Words, Cohn Matrices, and Cohn Tree
The Cohn matrix defined here differs from that in [1] in the arrangement of its entries. More precisely, the present definition is obtained from the one in [1] by swapping the diagonal entries.
@@
A proof of Theorem 2.11 can be found in [1]*Theorem 4.13. In [1], only the case $t\in[0,1]$ is treated, but the same argument applies to $t\in(1,\infty]$ as well.
@@ 2.3. Continued Fraction Matrix Decompositions of Cohn Matrices
Next, we define the strongly admissible sequence s(t)s(t) [^1] In [1], the term “strongly admissible sequence” is used for the bi-infinite sequence $^{\infty}s(t)^{\infty}$ in the sense of this paper, whereas here we use this name for its single block $s(t)$. . For $t=\frac{0}{1}$ or $t=\frac{1}{0}$, set $s(\frac{0}{1})=s(\frac{1}{0})=(1,1)$. For $t\in(0,\infty)$, we define $s(t)$ as follows:
@@
    If $t\in[1,\infty]$, then $c_{t}$ consists of letters $q$ and $r$, and $s(t)$ is obtained from $c_{t}$ by replacing each $q$ with “2,22,2” and each $r$ with “1,11,1”. In [1, 7], $s(t)$ is constructed from $c_{t}$ by using this replacement.
@@
By Remark 2.15 (2), $s(\frac{1}{t})$ is obtained from $s(t)$ by replacing the initial $2,2$ with $2,1,1$ and the final $1,1$ with $2$. The lemma follows from direct matrix computation (see also [8]*Corollary 24). ∎
@@ 2.4. Markov-Lagrange Spectrum and Markov Spectrum
We have the following theorem [^2] In [20], a more general formula is given for computing $\mathcal{L}(\alpha)$ for an arbitrary irrational number $\alpha$ using its continued fraction expansion and $\lim\sup$. By applying this formula to the case of quadratic irrationals, one obtains Theorem 2.20 stated in this section. For a detailed discussion of this reformulation, see [7]. :
Theorem 2.20 ([20]).
@@
The quadratic form $Q_{T}$ has rational coefficients. We have the following [^3] In [26], a more general formula is given for computing $\mathcal{M}(Q)$ for arbitrary irreducible quadratic form $Q$ using its continued fraction expansion and $\sup$. By applying this formula to the case of quadratic irrationals, one obtains Theorem 2.22 stated in this section. For a detailed discussion of this reformulation, see [22]*Appendix C.:
Theorem 2.22 ([26]).
@@ 3. Generalized Markov Equation and Generalized Markov Numbers
${((1,1),(4,2),(1,3))}$${((4,2),(21,1),(1,3))}$${((1,1),(17,3),(4,2))}$${((21,1),(121,2),(1,3))}$${((4,2),(457,3),(21,1))}$${((17,3),(373,1),(4,2))}$${((1,1),(81,2),(17,3))}$${((121,2),(703,1),(1,3))\cdots}$${((21,1),(15082,3),(121,2))\cdots}$${((457,3),(57121,2),(21,1))\cdots}$${((4,2),(10033,1),(457,3))\cdots}$${((373,1),(8185,3),(4,2))\cdots}$${((17,3),(38025,2),(373,1))\cdots}$${((81,2),(8227,1),(17,3))\cdots}$${((1,1),(386,3),(81,2))\cdots}$
@@
It follows from [2]*Proposition 3.8 and [18]*Theorem 1. ∎
@@
    If $b=\max\{a,c\}$, then $(a,b,c)=(1,1,1)$ ([18]*Lemma 4), therefore the set
@@
Let $(n^{\ast}_{\frac{1}{t}},i^{*}_{\frac{1}{t}})$ be a $(k_{1},k_{2},k_{3},\sigma^{\ast})$-GM number-position pair associated with $\frac{1}{t}$. Then, by Theorem 3.2, we have
@@
$u^{\ast}_{\frac{1}{t}}=n_{t}-u_{t}-k_{t}.$
The proof of Lemma 3.6 can be seen [16]*Proposition 7.22 (2) for the case where $k_{1}=k_{2}=k_{3}$. We can prove the general case in the same way. As in the case of Lemma 2.6, replace $v_{t}^{+}$ in [16] with $u^{*}_{\frac{1}{t}}$.
@@ 4.1. Generalized Cohn Matrices
In [17]*Lemma 4.5, Theorem 4.2 is treated for $k_{1}=k_{2}=k_{3}$ and $t\in[0,1]$. The general case is proved similarly.
@@ 4.2. Continued Fraction Matrix Decomposition of Generalized Cohn Matrices
To prove Theorem 4.6, we introduce the snake graph. Assume $a_{i}\in\mathbb{Z}_{\geq 1}$. First, recall the snake graph associated with a continued fraction $a_{1},\dots,a_{\ell}]$ according to [9]:
@@
In [9], the authors establish the following relation between a continued fraction and its associated snake graph:
Theorem 4.11 ([9]*Theorem 3.4).
@@
If $a_{1},\dots,a_{n}]=\frac{p_{n}}{q_{n}}$, then the corresponding product of continued fraction matrices is given by (see [7]*Section 2.1):
@@
On the other hand, [16]*Theorem 7.10 provides an expression for $C_{t}$ for $t\in(0,1)$. Adapting it to the notation of the present paper, we obtain the following:
@@
Note that when comparing Theorem 4.12 with [16]*Theorem 7.10 (2), the positions of the matrix entries differ due to the difference in the convention of Cohn matrices described in Remark 2.10.
@@
To prove Theorem 4.14, we need the snake graph calculus. The following two lemmas are essentially proved by [3]. In this paper, the statement is formulated in terms of the numerators of continued fractions, but by applying the Theorem 4.11 it can be rephrased as a statement about the number of perfect matchings.
Lemma 4.16 ([3]*Lemma 3).
@@
For the first statement of Lemma 4.17, the case where $k=1$ is proved in [3]*Lemma 5. We can prove other cases in the same way as [3]*Lemma 5. The second statement follows from Proposition 4.10. ∎
@@
The special case of Lemma 4.18 is proved by [16]*(7.2). Lemma 4.18 can be proved in the same way.
@@ 5. Generalized Markov Length and Generalized Markov Distance
Although the curve $\gamma_{AB}$ depends on whether the interior of the line segment from $A$ to $B$ is shifted to the left or right, the value of $d(A,B)$ is independent of this choice. This fact is proved by [5]*Lemma 8 for $k_{1}=k_{2}=k_{3}$, and we can apply the same argument to general cases.
@@
Theorem 5.7 ([5]*Proposition 2 and Lemma 9).
@@
In [5]*Lemma 8, only the case $k_{1}=k_{2}=k_{3}$ is treated, but the theorem applies in the same way even when these three values are distinct. In the case $k_{1}=k_{2}=k_{3}=0$, Theorem 5.7 is proved by [21]*Theorem 3.5.
@@ 6.1. Main theorems and their Proofs
$\ell(^{\infty}S|S^{\infty})=\frac{\sqrt{(\mathrm{tr}(CF_{S}))^{2}-(-1)^{n}4}}{(CF_{S})_{21}},$
@@
$F_{t}=\left\{\frac{\sqrt{(\mathrm{tr}(CF_{S}))^{2}-4}}{(CF_{S})_{21}}\ \middle|\ S=(a_{k}\dots,a_{k+n-1})\text{ for some $k$ }\right\},$
@@
By general theory of continued fraction matrix (see [8]*Appendix A), we have